# Blind Inverse Problem Solving Made Easy by Text-to-Image Latent Diffusion

### v1

@@ 3 Method
In this paper, we aim to tackle the problem of blind inverse problem solving defined in Equation 4. Our solution (Algorithm 1) has the following desiderata: (1) No additional training: our method should not require data collection or model training, (2) Adaptability to diverse image priors: the same model should apply to various image distributions, (3) Flexible operators: no assumptions about the operator’s functional form, such as linearity or task-specific update rules, should be necessary. To make this problem feasible, we assume access to open-sourced pre-trained models.
@@
This formulation allows us to decompose this problem into two parts: first obtaining a reliable estimation of the prior $p({\mathcal{A}}_{\phi})$ and then maximizing the measurement likelihood $p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})$. This makes diffusion-based framework the ideal approach as its posterior sampling process naturally separates these two stages (by Equation 2).
@@
1: /* Initialize latent with encoded measurement ${\bm{y}}$ SDEdit */
2:${\bm{z}}_{T_{s}}=\sqrt{\bar{\alpha}_{T_{s}}}{\mathbf{E}}(y)+\sqrt{1-\bar{\alpha}_{T_{s}}}\epsilon_{T_{s}}$ for $\epsilon_{T_{s}}\sim\mathcal{N}(0,I)$
@@
11:              /* Perform MPGD with the estimated ${\mathcal{A}}_{\hat{\phi}}$/
@@
14:         ${\bm{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}({\bm{z}}_{t},t)+\sigma_{t}\epsilon_{t}$ for $\epsilon_{t}\sim\mathcal{N}(0,I)$
15:         ${\bm{z}}_{t}=\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}{\bm{z}}_{t}+\sqrt{1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}\epsilon^{\prime}_{t}$ for $\epsilon^{\prime}_{t}\sim\mathcal{N}(0,I)$
16:    end for
17:     /* Periodically update $\hat{\phi}$ with ${\bm{z}}_{0|t}$ and Eq. 9 */
18:    if $t\equiv 0\pmod{5}$ then
19:         for $k=1,\dots,K$ do
20:             $\hat{\phi}=\text{Adam}(\hat{\phi},L_{\phi}({\bm{z}}_{0|t},\hat{\phi}))$
21:         end for
22:    end if
23:end for
24:return ${\bm{x}}_{0}={\mathbf{D}}({\bm{z}}_{0})$
3.1 Obtaining a better Prior
Most existing diffusion-based methods [8, 27, 34, 23, 1, 39] avoid modeling the joint prior and instead further decompose the goal distribution as $p({\bm{x}},{\mathcal{A}}_{\phi}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})p({\bm{x}})p({\mathcal{A}}_{\phi})$, under the assumption that the target data distribution $p({\bm{x}})$ independently from the operator distribution $p({\mathcal{A}}_{\phi})$. In this way, they can separately model the image and operator distribution by pre-trained diffusion models or fixed distributions and thus bypass the combinatorial complexity. However, is this truly a fair assumption?
Consider the task of deblurring and super-resolution. A low-resolution blurry measurement may correspond to a high-resolution blurry image (e.g. an out of focus DSLR photo) if the operator is a downsampler, or to a low-resolution clear image if the operator is a blur kernel. Similarly, different target image distribution can imply different operators: for example, the estimated blur kernel should differ if the target is a realistic photo versus an Impressionist painting, even when both share the same measurement.
These observations suggest that $p({\bm{x}})$ and $p({\mathcal{A}}_{\phi})$ should be correlated, and previous methods can make this independence assumption because they implicitly restrict the types of operations and target images they handle. In other words, while this independence assumption reduces modeling complexity, it also significantly limits the flexibility of the resulting algorithms. Consequently, most existing methods require additional training or parametrization of the operator prior and need to switch the image prior models when the target distribution changes.
Therefore, to build a unified training-free algorithm, we should directly model $p({\bm{x}},{\mathcal{A}}_{\phi})$, or in the case of diffusion models, $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$. On the first glance, modeling $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$ appears intractable. However, with large-scale pre-trained text-to-image diffusion models now widely available, we have a surprisingly simple solution:
In practice, the target data distribution (e.g. “a high-resolution photo”) and the forward operation (e.g. “Gaussian blur”) can often be effectively described in natural language. We can specify the target distribution by using its description as a positive prompt. Specifying the operator distribution is less direct, as these pre-trained models are designed to represent image distributions rather than operators. However, by leveraging classifier-free guidance (CFG) [15], we can implicitly define the operator distribution through negative prompts. For example, if the operator is Gaussian blur, an effective negative prompt can be “blurry image, low quality”. Formally, denote positive prompt as ${\bm{c}}_{+}$ and negative prompt as ${\bm{c}}_{-}$, we can approximate $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$ as
$\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})\approx\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{-})\\ +\gamma(\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{+})-\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{-}))$ (6)
where $\gamma>1$ is a weighting hyperparameter. This is equivalent to approximating $p({\bm{x}},{\mathcal{A}}_{\phi})$ as a distribution that is proportional to $p({\bm{x}}|{\bm{c}}_{-})^{1-\gamma}p({\bm{x}}|{\bm{c}}_{+})^{\gamma}$.
When using latent diffusion models parameterized by $\theta$, this corresponds to having the empirical noise prediction
This way, our method not only models this seemingly extremely intractable joint distribution in a remarkably simple way, it also enable wide range of applications with diverse target image and operator distributions by leveraging the knowledge in large-scale pre-trained text-to-image diffusion models like Stable Diffusion[31].
3.2 Obtaining a better Posterior
@@
2:     /* Initialize latent batch with encoded ${\bm{y}}$ SDEdit */
@@
5:     /* Use MPGD to obtain latent estimations $\{{\bm{z}}_{0|t}^{(i)}\}_{i=1}^{N}$/
@@
16:         $\hat{\phi}=\text{Adam}(\hat{\phi},\frac{1}{N}\sum_{i=1}^{N}L_{\phi}({\bm{z}}_{0}^{(i)},\hat{\phi}))$
@@
In addition to a strong prior, our final output image should also satisfy the measurement constraint.Given the problem setup in Equation 4, the measurements are subject to additive Gaussian noise ${\bm{n}}$, hence $\log p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})=-\frac{1}{2\sigma^{2}}\|{\bm{y}}-{\mathcal{A}}_{\phi}({\bm{x}})\|^{2}_{2}$. When ${\mathcal{A}}_{\phi}$ is known, we can use the MPGD update rule in Equation 3 to perform posterior sampling.
However, since ${\mathcal{A}}_{\phi}$ is not known, the true parameters $\phi$ is often approximated by another set of parameters $\hat{\phi}$. This approximation is usually addressed by one of the two strategies: an alternating optimization scheme that jointly approximates ${\bm{x}}$ and $\phi$, or obtaining a reliable $\hat{\phi}$ first then solving a non-blind inverse problem. The first approach is well-suited for training-free settings as it can be well integrated with diffusion-based methods, but it is often highly sensitive to $\hat{\phi}$ initialization and tuning. The second approach can perform well if a strong $\hat{\phi}$ is obtained, though it usually requires training and additional assumptions or restrictions.
We propose to combine the two strategies above: we first obtain a reliable initial $\hat{\phi}$, and then perform an alternating optimization to iteratively refine both the operator parameter and data estimations throughout the diffusion process.
@@
In some cases, initializing the operator parameters is straightforward – for example, for a slight blur, an identity function can be a reasonable stating point. Alternatively, reliable operator priors, such as the pre-trained prior in BlindDPS [8] or the Gaussian prior in GibbsDDRM [27] for linear operators, can also provide effective initializations for simple kernels. However, when the operator requires complex functions like neural networks, obtaining good initializations becomes more challenging.
To address this, we propose a new algorithm for general operator initialization by drawing further inspiration from alternating optimization algorithms. Note that since the goal here is only to initialize $\hat{\phi}$, the quality of intermediate ${\bm{x}}$ estimates is less critical as long as they provide a useful signal. Unlike other diffusion-based methods that alternate optimization targets at each diffusion step, we use SDEdit [26] and MPGD [14] with very few diffusion steps to quickly obtain a batch of ${\bm{x}}$ estimates. We then perform maximum likelihood estimation (MLE) on these estimates to update $\hat{\phi}$ using Adam optimizer. As detailed in Algorithm 2, repeating this process can leverage the diffusion prior to quickly converge to a reliable starting point for $\hat{\phi}$.
Notice that we only assume that $\phi$ can be approximated by differentiable functions, allowing ${\mathcal{A}}_{\hat{\phi}}$ to be parameterized by general model families such as neural networks. Moreover, as mentioned earlier, any well-performing operator priors can be seamlessly integrated into our framework.
@@
After initializing $\hat{\phi}$, we then perform another alternating optimization scheme to refine both the operator and the data. In particular, we alternate between updating ${\bm{z}}_{0|t}$ using MPGD guidance with ${\mathcal{A}}_{\hat{\phi}}$ fixed and updating the MLE estimation of $\hat{\phi}$ using Adam with ${\bm{z}}_{0|t}$ fixed throughout the diffusion process.
Since unlike GibbsDDRM, which uses Langevin dynamics to update the operator, we solve for a local optimum. Therefore, it’s reasonable not to update the operator too frequently. Empirically, we find that periodic updates to the operator in combination of the time-traveling technique [25, 42] yield the best results
In addition, since MPGD supports any differentiable loss function, we can incorporate techniques like regularization to further improve the visual quality. In practice, we use
@@
By combining large-scale language-conditioned diffusion priors, effective operator initializations, and the iterative refinement process described above, we introduce LADiBI (Algorithm 1), a new training-free algorithm for blind inverse problem solving that supports diverse target image distributions and flexible operators.
@@ 4.1 Experimental Setup
We empirically verify the performance of our method with two linear deblurring tasks, Gaussian deblurring and motion deblurring, and a non-linear restoration task, JPEG decompression. Here we provide the details of our experiments.
Datasets
We conduct quantitative comparisons on the datasets FFHQ 256 $\times$ 256 [18] and AFHQ-dog 256 $\times$ 256 [6]. We use a size of 1000 images for FFHQ and 500 for AFHQ, following the setting in Murata et al. [27].
@@
We compare our method against other state-of-the-art approaches as baselines. Specifically, we choose Pan-l0 [28] and Pan-DCP [29] as the optimization-based method, SelfDeblur [30] as the self-supervised approach, PRNet [44] and DeblurGANv2 [22] as the supervised baselines, and BlindDPS [8] and GibbsDDRM [27] as the diffusion-based methods. All baselines are experiments using their open-sourced code and pre-trained models, whose details are provided in the appendix.
@@
We conduct all experiments an NVIDIA A6000 GPU. We use Stable Diffusion 1.4 [31], DDIM $200$-step noise scheduling, and $T_{s}=150,M=4$ for all tasks. We set positive prompts as “a clear headshot of a person” and “a clear headshot of an animal” for FFHQ and AFHQ experiments respectively. We adjust negative prompts according to each individual task, and the details will be provided in the following sections.
Evaluation Metrics
We evaluate our method and the baselines with three quantitative metrics: LPIPS [46], PSNR and KID [4]. Each metric has its own significance and reveals a different set of characteristics about performance. In particular, PSNR directly measures the pixel-by-pixel accuracy of the reconstructed image compared to the original, while LPIPS measures how similar two images are in a way that reflects human perception. Finally, KID evaluates the image fidelity on a distribution level.
4.2 Blind Linear Deblurring
We first test our method against the baselines on linear deblurring, which is the design space of most baselines. We evaluate all methods on two types of blur kernels: Gaussian blur and motion blur.
Setup
To generate measurements, we apply randomly generated motion blur kernels with intensity 0.5 and Gaussian blur kernels with standard deviation 3.0 respectively. The final measurements are derived by applying a pixel-wise Gaussian noise with $\sigma=0.02$. We use a $61\times 61$ convolutional matrix as $\hat{\phi}$ and we initialize it as a Gaussian kernel of intensity 6.0 following Murata et al. [27]. We adjust the negative prompts to the potential negative image artifacts induced by the degradation operation, such as “lowres, blurry, DSLR effect, poorly drawn face, deformed, bad proportions, …” for Gaussian blur, “shaken image, morbid, mutilated, text in image, disfigured, …” for motion blur. Full prompts are included in the appendix.
Results
@@
Table 2 and Figure 2 show results on blind linear deblurring. Our method matches the performance the state-of-the-art method GibbsDDRM, which is explicitly designed to solving linear problems using SVDs. In fact, GibbsDDRM’s highly specialized design makes it so sensitive that even small discrepancies (e.g. kernel padding) between their modeling assumption and the ground truth operator can result in significant performance degradation. Moreover, although BlindDPS and GibbsDDRM use diffusion models that are trained on the exact data distributions tested, our method can still match or outperform them with the large-scale pre-trained model. In addition, even though supervised methods can obtain higher PSNR, our method produces the fewest artifacts, which is also validated by the LPIPS and KID scores. Overall, our method offers a consistently competitive performance on linear tasks, even though our method is designed for more general applications.

### v2

@@ 3 Method
In this paper, we aim to tackle the problem of blind inverse problem solving defined in Equation 4. Our solution has the following desiderata: (1) No additional training: it should not require data collection or model training, (2) Adaptability to diverse image priors: the same model should apply to various image distributions, (3) Flexible operators: no assumptions about the operator’s functional form, such as linearity or task-specific update rules, should be necessary. To make this problem feasible, we assume access to open-sourced pre-trained models.
@@
This formulation allows us to decompose this problem into two parts: obtaining sampling to maximizes the measurement likelihood $p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})$. This makes diffusion-based framework the ideal approach as its posterior sampling process naturally separates these two stages.
Figure 2: A schematic overview of LADiBI (Algorithm 1).
3.1 Obtaining the prior score
As established in Equation 2, diffusion-based approaches to inverse problems require access to the score of the prior distribution. For blind inverse problems, this extends to estimating the score of the joint distribution $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$, which presents significant challenges beyond standard image priors. While strong pre-trained diffusion models usually exist for target image distributions $p({\bm{x}})$, appropriate models or even training data for degradation operators ${\mathcal{A}}_{\phi}$ are largely unavailable, making estimations of the joint distribution $p({\bm{x}},{\mathcal{A}}_{\phi})$ particularly difficult.
Therefore, most existing diffusion-based methods further decompose the joint prior into independent distributions $p({\bm{x}},{\mathcal{A}}_{\phi})\approx p({\bm{x}})p({\mathcal{A}}_{\phi})$ and impose restrictive assumptions, such as linearity constraints, on the operator (Chung et al., 2023a; Murata et al., 2023; Sanghvi et al., 2023; Laroche et al., 2024; WeiminBai et al., 2025; Tu et al., 2024). To estimate the operator distribution, they either rely on simulated training data tailored to specific operator forms or perform constrained optimization under fixed operator distributions. In addition, these methods typically require different pre-trained diffusion models for different image domains. For example, deblurring human faces versus animal faces necessitate different prior models. While effective in constrained settings, their reliance on domain-specific models and strong operator assumptions fundamentally limits their generalizability, particularly violating the second and third desiderata.
Motivated by these limitations, we introduce an alternative approach inspired by the following key insights. First, many image restoration tasks can be intuitively described using natural language. In addition, large pre-trained text-to-image diffusion models like Stable Diffusion (Rombach et al., 2021) already capture rich distributions of both target images and common degradation artifacts described by these prompts. Based on these insights, we propose to approximate the joint prior’s score using large-scale pre-trained text-to-image diffusion models through classifier-free guidance (Ho and Salimans, 2021). In particular, we encode the desired target image characteristics using the positive prompts, and specify the unwanted degradation artifacts via a negative prompts. For example, when restoring a human face degraded by Gaussian blur, a suitable positive prompt can be “a clear headshot of a person”, and the corresponding negative prompt could be “blurry, low-quality image”.
Formally, denoting the positive prompt as $\bm{c}$$_{+}$ and the negative prompt as $\bm{c}$$_{-}$, we can approximate $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})$ as
$\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t},{\mathcal{A}}_{\phi})\approx\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{-})+\gamma(\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{+})-\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{c}}_{-}))$ (6)
where $\gamma>1$ is a weighting hyperparameter. When using latent diffusion models parameterized by $\theta$, this translates to the empirical noise prediction
This straightforward approximation provides access to the otherwise intractable joint prior score, and enables versatile applications across diverse image and operator distributions. By leveraging the knowledge encoded in large pre-trained text-to-image models, our method bypasses the need for task-specific training, operator restriction or model re-selection.
While this approach may appear to compromise the true "blindness", we emphasize that all blind inverse problem methods require some prior over the transformation space. Notably, textual descriptions of the target image and degradation artifacts are naturally advantageous, since users can directly infer the appropriate prompts from the measurements, which is a significantly simpler process than previous methods. These prompts provides an intuitive interface to encode diverse characteristics of both the desired outputs and the unwanted artifacts without training or handcrafted functions.
@@ 3.1 Obtaining the prior score
1: /* Initialize latent with encoded measurement $\bm{y}$} & SDEdit */
[…]
@@
11:      /* Perform MPGD with the estimated $^{\mathcal{A}}_{\hat{\phi}}$ */
@@
14:   $\epsilon_{t},\epsilon^{\prime}_{t}\sim\mathcal{N}(0,I)$
15:   ${\bm{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}({\bm{z}}_{t},t)+\sigma_{t}\epsilon_{t}$
16:   ${\bm{z}}_{t}=\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}{\bm{z}}_{t-1}+\sqrt{1-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}\epsilon^{\prime}_{t}$
17:  end for
18:   /* Periodically update $\hat{\phi}$ with ${\bm{z}}_{0|t}$ and Eq. 9 */
19:  if $t\equiv 0\pmod{5}$ then
20:   for $k=1,\dots,K$ do
21:     $)\hat{\phi}=\text{Adam}(\hat{\phi},L_{\phi}({\bm{z}}_{0|t},\hat{\phi}$))
22:   end for
23:  end if
24:end for
25:return ${\bm{x}}_{0}={\mathbf{D}}({\bm{z}}_{0})$
@@
2:   /* Initialize latent batch with encoded $\bm{y}$} & SDEdit */
@@
5:   /* Use MPGD to obtain latent estimations $\{{\bm{z}}_{0|t}^{(i)}\}_{i=1}^{N}$ */
@@
16:   $)\hat{\phi}=\text{Adam}(\hat{\phi},\frac{1}{N}\sum_{i=1}^{N}L_{\phi}({\bm{z}}_{0}^{(i)},\hat{\phi}$))
@@
3.2 Sampling from the posterior
In addition to a strong prior score, our output image should also satisfy the measurement constraint. Given the problem setup in Equation 4, the measurements are subject to additive Gaussian noise $\bm{n}$}, hence $\log p({\bm{y}}|{\bm{x}},{\mathcal{A}}_{\phi})=-\frac{1}{2\sigma^{2}}\|{\bm{y}}-{\mathcal{A}}_{\phi}({\bm{x}})\|^{2}_{2}$. When ${\mathcal{A}}_{\phi}$ is known, we can use the MPGD update rule in Equation 3 for posterior sampling.
However, since ${\mathcal{A}}_{\phi}$ is unknown, the true parameters $\phi$ is often approximated by another set of parameters $\hat{\phi}$. This approximation is usually addressed by one of the two strategies: an alternating optimization scheme that jointly approximates $\bm{x}$} and $\phi$, or obtaining a reliable $\hat{\phi}$ first then solving a non-blind inverse problem. The first approach is well-suited for training-free settings, but it is often highly sensitive to $\hat{\phi}$ initialization and tuning. The second approach can perform well if a strong $\hat{\phi}$ is obtained, though it usually requires training and additional restrictions. We propose a hybrid strategy: we first obtain a reliable initial $\hat{\phi}$, and then perform an alternating optimization to iteratively refine both the operator parameter and data estimations throughout the diffusion process.
@@ 3.2 Sampling from the posterior
Sometimes initializing the operator parameters is straightforward – for example, for a slight blur, an identity function can be a reasonable starting point. Alternatively, reliable operator priors, such as the pre-trained prior in BlindDPS or the Gaussian prior in GibbsDDRM for linear operators, can also provide effective initializations for simple kernels. However, for complex operator estimator like neural networks, obtaining good initializations becomes more challenging.
To address this, we propose a new algorithm for general operator initialization. Since the goal here is only to initialize $\hat{\phi}$, the quality of intermediate $\bm{x}$} estimates is less critical as long as they provide useful signals. Unlike other diffusion-based methods that alternate optimization targets at each diffusion step, we use SDEdit (Meng et al., 2022) and MPGD (He et al., 2024) with very few diffusion steps to quickly obtain a batch of $\bm{x}$} estimates. We then perform maximum likelihood estimation (MLE) on these estimates to update $\hat{\phi}$ using Adam optimizer. As detailed in Algorithm 2, repeating this process can leverage the diffusion prior to quickly converge to a reliable starting point for $\hat{\phi}$.
Notice that we only assume that $\phi$ can be approximated by differentiable functions, allowing $^{\mathcal{A}}_{\hat{\phi}}$ to be parameterized by general model families such as neural networks. Moreover, as mentioned earlier, any well-performing operator priors can be seamlessly integrated into our framework.
@@
After $\hat{\phi}$ initialization, we perform another alternating optimization to refine both the operator and the image. Throughout the diffusion process, we alternate between updating ${\bm{z}}_{0|t}$ using MPGD with $^{\mathcal{A}}_{\hat{\phi}}$ fixed and updating the MLE estimation of $\hat{\phi}$ using Adam with ${\bm{z}}_{0|t}$ fixed.
Since unlike GibbsDDRM, which uses Langevin dynamics to update the operator, we solve for a local optimum. Therefore, it’s reasonable not to update the operator too frequently. Empirically, we find that periodic updates to the operator combined with time-traveling (Lugmayr et al., 2022; Yu et al., 2023) yield the best results. In addition, since MPGD supports any differentiable loss function, we can incorporate regularization to further improve the visual quality. In practice, we use
@@
Combining large-scale language-conditioned diffusion priors, effective operator initializations and the iterative refinement process described above, we introduce LADiBI, a new training-free algorithm for blind inverse problem solving that supports diverse target image distributions and flexible operators.
@@ 4.1 Experimental Setup
We empirically verify the performance of our method with two linear deblurring tasks, Gaussian deblurring and motion deblurring, and a non-linear restoration task, JPEG decompression.
Setup
We conduct quantitative comparisons on FFHQ 256 $\times$ 256 (Karras et al., 2019) and AFHQ-dog 256 $\times$ 256 (Choi et al., 2020). Following Murata et al. (2023), we use 1000 images for FFHQ and 500 images for AFHQ. We use three quantitative metrics: LPIPS (Zhang et al., 2018), PSNR and KID (Bińkowski et al., 2018). PSNR directly measures the pixel-by-pixel accuracy of the reconstructed image compared to the original, while LPIPS measures how similar two images are in a way that reflects human perception. Finally, KID evaluates the image fidelity on a distribution level.
@@
We compare our method against other state-of-the-art approaches as baselines. Specifically, we choose Pan-l0 (Pan et al., 2017) and Pan-DCP (Pan et al., 2018) as the optimization-based method, SelfDeblur (Ren et al., 2020) as the self-supervised approach, PRNet (Zamir et al., 2021) and DeblurGANv2 (Kupyn et al., 2019) as the supervised baselines, and BlindDPS (Chung et al., 2023a), BIRD (Chihaoui et al., 2024) and GibbsDDRM (Murata et al., 2023) as the diffusion-based methods. All baselines are experiments using their open-sourced code and pre-trained models.
@@
We conduct all experiments on an NVIDIA A6000 GPU. We use Stable Diffusion 1.4 (Rombach et al., 2021), DDIM $200$-step, and $T_{s}=150,M=4$ for all tasks. We set positive prompts as “a clear headshot of a person" and “a clear headshot of an animal" for FFHQ and AFHQ experiments respectively. We adjust negative prompts according to each individual task and the details are be provided in the Appendix A.
@@
4.2 Blind Linear Deblurring
We first conduct experiments on linear deblurring, which is the design space of most baselines. We evaluate all methods on two blurring kernels: Gaussian blur and motion blur. We apply random motion blur kernels with intensity $0.5$ and Gaussian blur kernels with standard deviation $3$. Measurements are derived by applying a pixel-wise Gaussian noise with $\sigma=0.02$. We use a $61\times 61$ convolutional matrix as $\hat{\phi}$ and initialize it as a Gaussian kernel of intensity $6.0$ following Murata et al. (2023)
As shown in Table 2 and Figure 3, our method matches the performance of the state-of-the-art method GibbsDDRM, which is explicitly designed to solve linear problems using SVDs. In fact, GibbsDDRM’s highly specialized design makes it so sensitive that even small discrepancies (e.g. kernel padding) between their modeling assumption and the ground truth operator can result in significant performance degradation. Moreover, although BlindDPS and GibbsDDRM use diffusion models that are trained on the exact data distributions tested, our method can still match or outperform them with the large-scale pre-trained model. While supervised methods obtain higher PSNR, our method produces the fewest artifacts, validated by LPIPS and KID scores. Overall, our method offers a competitive performance on linear tasks, even though it is designed for more general applications.